OpenStudy (anonymous):

Which inequality models this problem? The length of a rectangle is four times its width. If the perimeter is at most 130 centimeters, what is the greatest possible value for the width? A. 2w + 2 • (4w) < 130 B. 2w + 2 • (4w) > 130 C. 2w + 2 • (4w) ≤ 130 D. 2w + 2 • (4w) ≥ 130

OpenStudy (texaschic101):

P = 2L + 2W L = 4W P <= 130 2(4W) + 2W <= 130 or 2W + 2 * 4W <= 130 (C)